I really stink at written problems. Could you check to see if my answer is corret. If not, help.
The base of the triangle is 4 inches longer than the length of one of the other sides. Use a system to equations to find the length of each side of the triangle.
x in
x in
Perimeter= 46 in
y in
My answer is x=14 and Y =18

Answer by checkley75(3666)

your identification of the triangle is x,x&y which says it is an isosceles triangle then
x+x+(x+4)=46
3x+4=46
3x=46-4
3x=42
x=42/3
x=14 for the equal sides & 14+4=18 for the base.proof
14+14+18=46
46=46